OpenStudy (anonymous):

what is the equation of the line in standard form that passes through the point (1,5)and is parallel to the line y=-3x-3?

OpenStudy (aroub):

parallel= same slope

OpenStudy (anonymous):

yup so y=-3x+b 5=-3(1)+b 5=-3+b 8=b y=-3x+8

OpenStudy (aroub):

y=mx+b , m= -3 , (1,5) 5=-3(1)+b 5=-3+b 8=b so the equation is y=-3x+8

OpenStudy (aroub):

lana!! xx thank yoouuu

OpenStudy (lalaly):

<3

OpenStudy (anonymous):

thanks

OpenStudy (aroub):

anytime!! :)